wiki:UC20191105

Version 1 (modified by trac, 11 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 11/05/19

Attending: Kevin Schlaufman (JHU), Ben Williams (UW), Mukremin Kilic (OU), John Bally (CU), Kal Kadlec (UW), Nancy Chanover (NMSU), Derek Buzasi (FGCU), Eric Bellm (UW), Russet McMillan (APO), Joanne Hughes Clark (Seattle U), Mark Klaene (APO)


User feedback and comments from institutional representatives:

  • Georgia State - no report
  • Washington - nothing to report
  • UVa - nothing to report (via email)
  • Seattle - nothing to report
  • JHU - nothing to report
  • Wyoming - no report
  • FGCU - nothing to report
  • Oklahoma - nothing to report
  • NAPG - no report
  • Colorado - nothing to report
  • NMSU - nothing to report (via email)

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 10/03/2019 - 10/28/2019

0) Overview

We had minimal problems this period other than a few oddities that could be one-offs. Good progress was made on the ARCSAT guider. The site received a lightning strike that caused some communications issues with ARCSAT and also blew two holes in the water line along the road coming from Sunspot. This took several days to locate and repair, resulting in a water outage throughout that period.

1) Telescope

Nominal operations with minimal problems. Some tracking issues occurred, which we think were primarily due to seasonal temperature changes. Bill and Shane are looking at how to tweak the tuning. The problem is not unusual, nor did it cause significant loss of time.

2) Instruments

DIS: The blue camera contamination is still bad and there are no plans to service it. The red camera is getting worse but is not considered bad yet. Bill made a change to the ion pump configuration and as a result the red ion pump has settled down.

TripleSpec is operational. A pattern noise was noted at a low level on some images and UVa is looking into it.

Agile is operational.

Echelle is operational. The inter-order light In blue has improved. One odd readout error occurred and we are monitoring it. Ecam windowing with binning is still an a issue and is being looked at.

NIC-FPS is operational.

ARCTIC is operational.

3) ARCSAT

Operating nominally. The new guider was configured and commissioned and seems to be working well.


Additional telescope and instrument discussion:

No additional discussion took place beyond summarizing Mark's monthly report.


Q4 3.5m and ARCSAT scheduling:

Only two unassigned half-nights remain in Q4; both have been requested but we have not yet committed to those requests. Anyone with particularly urgent or compelling requests can submit them, but soon we will have no unassigned time remaining in the schedule. This is assuming that the KOSMOS commissioning will take place as scheduled (during the week of December 9), so if proposers have a need for extra time and they have substantial schedule flexibility in December, they can indicate their availability in the event that the KOSMOS commissioning dates change.


Q4 ARCSAT scheduling:

There is a little open time left in the Q4 schedule but not a lot. We have recently seen a fair number of requests for changes in filters. When submitting requests for time, proposers should be sure to specify both instrument and filter choices (or indicate flexibility in these areas).


Q1 3.5m scheduling:

The institutional allocations were all sent out yesterday, so everyone should be hearing from their institutional schedulers soon if they have not already. All requests are due to Ben and Russet on November 27, 2019.


NA2 Baffle Project:

The pre-fabrication review was completed and the UVa team followed up with some additional information that quantifies the expected scattered light on the ARCTIC detector as a result of the new (larger) NA2 baffle and the expected impact on guider performance. These data are available for anyone who is interested; please contact Nancy for more details. Having passed the pre-fab review, the NA2 baffle is now in fabrication and is expected to be completed by the end of CY19, with commissioning expected to be in early Q1 2020.


KOSMOS Update:

The instrument aside from the dewar is pretty much complete. The UW team is doing a lot of programming of motors and the like but everything is assembled. A recent hiccup is that last week they found that the CCD was having voltage issues. The detector was sent to Imaging Technology Laboratory at University of Arizona for inspection and hopefully repair, but it is not clear when that will be done. The UW team is exploring the possibility of using the engineering grade detector for the purposes of finishing their lab tests but it unclear whether that will be of value. The current commissioning schedule is still what is being targeted, but since we are unsure of the scale of repairs required for the detector, it is not yet known whether the commissioning schedule will have to slip. We should find out more in the next week or so.


Status Update on Near-Term Action Items from Strategic Planning Meeting

  • DAWG: The DAWG implemented a survey to assess the tools, pipelines, and software packages that people are using to reduce their APO data. Now that the deadline for responses has passed, they will likely have a telecon in the near future to discuss the survey results.
  • Getting narrow-band filters for ARCTIC: Nancy will discuss this with the management team and make a decision so she can get them ordered before the end of CY19.
  • Science communications intern: Nancy interviewed several candidates and will be making an offer today.
  • AAS presence: Nancy reserved an exhibitor booth, and needs to know who from ARC and leasing institutions is a) going to the AAS meeting in January and b) willing to help out by spending some time at the booth talking about APO to passers-by. After polling the UC reps on the call and based on other input Nancy received, the following people are going and willing to spend time at the booth: Nancy, Ben, Mike Crenshaw, Karen Kinemuchi, Nickalas Reynolds. UC reps: please email your department members and follow up on this to see whether there are others in your departments who can help out at the booth.
  • Student training videos: Nancy will contact trip leaders to see what might be possible.

Report from ARC Board of Governors Meeting

Nancy recapped her presentations to the Board (and will make them available on the wiki). She summarized some of the main points, and discussed the following key take-away points from the meeting:

  • There is a growing sense of urgency that we need to improve our instrumentation. Plans to address the aging DIS are in place (KOSMOS as a stop-gap measure, followed by the new spectrograph), but we also need to address the aging echelle. Nancy is following up on some leads for new echelle options that were discussed at the Strategic Planning meeting, and is assembling a small group of active echelle users who might be willing to help advise on this process. This is not expected to be a major time commitment, rather, a standing group who would help Nancy evaluate options. She asked for suggestions of people who could help with this and the following people were proposed or volunteered: Kevin Schlaufman (JHU), Derek Buzasi (FGCU), Yuta Notsu (UW -> CU), George Wallerstein and Charli Sakari (UW). UC reps: please poll your users, particularly those with a vested interest in an echelle replacement, and ask for volunteers to serve on this advisory group.
  • We requested a budget increase for the 3.5m to keep up with staff salary increases (which are a good thing!!) and the rising costs of facility maintenance and repairs. The Board supported this request, which sends a strong signal that the Board is committed to retaining our excellent staff and keeping the facility strong, and that they support the long-term health of the 3.5m.
  • Mark has announced his retirement; he plans to retire in Summer 2020. The job ad for his successor is now posted (e.g. see https://jobregister.aas.org/ad/5c03e128) and the search committee is eager to identify a broad and excellent applicant pool. If anyone knows of any potential candidates please notify Nancy (as a member of the search committee). However, Mark is truly irreplaceable! The Board acknowledged their appreciation of his tireless work and dedication to APO for the past several decades, and Nancy reiterated her thanks and appreciation of Mark on today's call. The UC will be updated on the search for the next Director of Operations at APO as things develop.
  • Nancy showed the Board a video that she had made commemorating the 25th anniversary of APO. She will provide the link as soon as it gets posted to a generally accessible site.

ACTION ITEMS:

Open action items from previous meetings:

  • Nancy will contact trip leaders bringing student groups to APO in Q4 to discuss training video ideas. STATUS: OPEN.
  • UC reps should poll their users to see who is going to the AAS meeting and would be willing to help out with booth presence. STATUS: OPEN.
  • UC reps should request science highlights from their users and have them sent to Nancy no later than October 15. STATUS: CLOSED.

New action items from this meeting:

  • UC reps should poll their users, particularly those with a vested interest in an echelle replacement, and ask for volunteers to serve on the echelle replacement advisory group.

All Other Business

None.


Next meeting:

The next Users Committee telecon will be on Monday December 2 at 10:30 am MST.


Note: See TracWiki for help on using the wiki.